Shift Summary for eve shift, Sept 23, 1999

Patricia McBride (mcbride@fnpsph)
Fri, 24 Sep 1999 00:16:19 -0500 (CDT)

Shift Summary for eve shift, Sept 23, 1999

On Shift: P. McBride, J. Graham

Hours of beam available at all intensities: 5.5
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 1.5
Hours of special runs (laser scans, muon runs...): .75

Accounting of lost time: recovery from shutdown; access to
replace pipeline power supply

========================================================================
Runs ended this shift:
14694 LASER
14697 KTEV_PED
14699 MUON_4E12 0.5 M events
14700 BEAM_799 short run, stopped because ptkick was low
14701 BEAM_799 1.0 M stopped run when pipeline power supply failed

========================================================================
Accesses during shift:

BobT and JimG made an access to replace a pipeline power supply

========================================================================
Overview of shift:

Beam returns in record time
MCR says that they are having trouble with loses in the Tevatron and
can only run 1 turn for the moment. They are investigating. In the
meanwhile we are taking data at 4e12.

Access to replace pipeline power supply.

After the access, the beam returns at 8e12.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14699 1 18:36 Start muon run

3 muon run ended with a rad trip in NM3. Trip might
be real. MCR plans to monitor the situation.
In the event of another rad trip, they will contact
Mike Gerardi and diagnose the problem further.

14700 19:45 Negative polarity BEAM_799
ERRORs: ptkick not correct, e/p low by 1%;
end the run, reset magnet to zero and
go back to positive polarity

14701 1 19:37 Similar problems with the magnet. E/p still low by 1%
We decide to continue so we page Ashkan to ask him
to stop the alarm. Intensity at 4e12.

~60 Pipeline power supply fails; BobT and JimG go on
access
We stopped the run and updated the T0s. (DC3Y window
wire
voltage was changed to 2200 V today).

14702 1 23:55 Jim and Bob return from access having replaced the
power supply
call MCR, beam returns at 8e12

========================================================================
Follow up Notes on Previous Shift

Jason changed HV for some CSI channels. (See Logbook XXI, p284)

========================================================================
Items to Follow up on Next Shift

Remember to flip the magnet to negative at the end of the next run.